ESMOS STATUS MEETING MINUTES
The ESMOS Status Meeting was conducted by K. Dolan at 1:00 PM March 26, 2002. The meeting was held in Room N202 in Building 32.
General Comments and Related Discussion:
Cannot get the MODIS instrument up since Terra went into safe mode. There is a reset problem; also have not been able to load patches yet.
The batteries on Aqua may be okay. Battery testing has delayed fueling. If the batteries fail during testing, launch will be delayed until the batteries are replaced.
Monthly status review inputs are due.
The official Delta ORR for UARS is April 4th.
Discussion of ESMOS Top Ten Issues
The Mission System top ten issues were discussed by K. Dolan.
#1 FY02 budget - SOMO backing out of funding networks, institutional support and technology programs: Meetings are being held. Working on FY03 budget.
#2 Readiness of Ground Network to support Aqua: Next GN test is scheduled for April 8th through the 10th. Approximately 95% of the data has been collected during the GN tests.
#3 NACs: Almost completed including those from TRW/Redondo Beach.
#4 Getting a contract in place to provide anomaly support from LMMS, SBRS, and L3COM: There are still some issues with labor rates; an update is expected.
#5 End of Mission planning for ERBS and UARS: Meetings were held with W. Townsend, Code Q, and Code Y. Some action items were assigned. For ERBS, we need to provide probabilities of creating debris in orbit; need a response in one month.
#6 NCR: All updates have been made.
#7 USGS has not provided money to cover FY02 Landsat-7 network costs. NASA has not budgeted for this cost: Nothing new to report.
#8 DSMC transition NCC and WOTIS operations are moving to WSC during 2002: Nothing new to report.
#9 System vulnerabilities: Nothing new to report.
#10 Aqua Transition: No updates; final transition plans is not due yet.
Comments by K. Dolan and Related Discussion:
Meetings listed on the web site calendar and were reviewed.
Comments by B. Guit and Related Discussion:
The UARS re-engineering effort was very successful; the system was completed ahead of schedule and under budget.
Mid-term report was given on science traceability mission.
Comments by R. Casey and Related Discussion:
Terra inclination maneuver 5 is being reviewed as part of the preparation for inclination maneuver 6.
Comments by C. Adams and Related Discussion:
Aqua updates for the WOTIS ICD are being reviewed.
A revision to the Aura MSRD will be out this week; Raytheon was given an advanced copy.
Comments by A. Jansen and Related Discussion:
Howard University students (about 25) will be here this March 29th for a walking tour as part of the outreach program.
Comments by A. Kelly and Related Discussion:
Aura project personnel are at a meeting at JPL. The project is getting ready for SDVF testing with each of the instruments. This testing is scheduled to start in late April and continue until mid-May. There should some instrument data collected.
Comments by L. Kurzmiller and Related Discussion:
TRMM data that was lost during data drop out at the 1513Z event was recovered on the following event at 1602Z. TRR 24263 and event report 259 were prepared.
Comments by J. Wallace and Related Discussion:
Need to fix some minor non-launch critical load-center switching relay problems prior to Aqua launch. These problems were found during the last load-center relay testing.
Comments by R. Baldwin and Related Discussion:
April 9th will be the soft freeze date for infrastructure activities based on an April 18th launch date. Any infrastructure activities that could impact Aqua mission critical areas will be halted. The Building 32 infrastructure will be placed in a hard freeze two days prior to launch. Slips will be linear.
Comments by B. Seaton and Related Discussion:
Current plans are to complete the migration of ASTER devices by the end of the summer.
EDOS operational readiness testing is in progress.
EDOS Release C4.2ae is being installed.
Carlos Gomez is investigating the back jumps in time in ASTER data; the jumps appear to occur in spikes. Andy Jansen stated that Dan Linebarger or Eric Moyer should be contacted for analysis support from the Terra FOT.